Donna Kirkpatrick formed Racing Cars Corporation early this century with an investment of $5 million in cash, for which she received $4 million in common stock and $1 million in bonds bearing interest of 9% with maturity in 2015. In 2007, Donna lent another $1 million on a personal note with 2% interest. This year Racing Cars becomes insolvent after Donna tries to take the first turn at an unauthorized speed of 250 mph, destroying the last car owned by Racing. Sponsors pulled all of their support, and Donna had to send pitiful termination letters to the pit crew and staff. Racing Cars Corporation has to file bankruptcy. During the last two years, Donna was paid an annual salary of $10 million. Donna wants you to consider the above information and explain to her the possible tax positions of Racing in the last year of business. She also wants to know her personal tax situation considering the worthless stock she now uses as floor mats in her new pickup. How would you handle this tax situation?
Answer #1

Investments=$5million

Stock=$4 million
Bond=$1million
Interest on bonds=9 %=( 9%=9/100)*$ 1million = 90,000 creditors = (102%/100)*1 million =1.02 million

As per IRS (Internal Revenue Service), the requirement to file a return for a bankruptcy estate applies if gross income is at least $12,000. As it is mentioned in the question that the last racing car was destroyed because of high speed.The company became insolvent. So the gross income is 0 (zero). It means that company does not have to pay anything as it has paid somewhat to creditors based on what it had.

Once Donna files for bankruptcy the government will stop from collecting taxes that Donna owes until the bankruptcy is discharged. If anything is due a refund, it is considered as an asset. That the money can be used to pay the creditors. The salary which Donna has got (10$ million) can be used to pay creditors.

REFERENCES: IN APA FORMAT

Perritano, J., 2014. How does bankruptcy affect your tax return? HowStuffWorks. Available at: https://money.howstuffworks.com/personal-finance/personal-income-taxes/bankruptcy-tax-return.htm [Accessed January 24, 2020].

Asatryan, M., 2015. Tax Debts in Chapter 7 Bankruptcy. www.nolo.com. Available at: https://www.nolo.com/legal-encyclopedia/tax-debt-chapter-7-bankruptcy.html [Accessed January 24, 2020].